OpenStudy (countrygirl1431):

Solve the following equation. 8 + x = 13 + 2x

OpenStudy (anonymous):

To solve this, we have to isolate the variable, or get it alone on one side of the equation. Since subtraction is the inverse operation for addition, we can subtract things from both sides to accomplish this.\[8+x-x=13+2x-x\] \[8=13+x\] \[8-13=13-13+x\] \[-5=x\]

OpenStudy (countrygirl1431):

3x + 1 + 5x = 7 +15 + 7x i tried this problem so many ways and i could never get it to come out right

OpenStudy (anonymous):

this is how i did the equation. Step 1. pull out like terms / factors so your problem would look like this -5-x=-1*(x+5)=0 Step 2. solve -1*(x+5)=0 part1. -1=0 part2. solve the single variable equation solve x+5=0 then subtract 5 from both sides and that should give you your answer